Given the demand function

Pd = 35 − Q2,

and the supply function

Ps = 3 + Q2.

Calculate the producer surplus for the supply function.


1
Expert's answer
2021-07-14T13:59:18-0400

At equilibrium

"PD=PS"

"(35-Q^2)=(3+Q^2)\\
32 =2Q^2\\
Q=4"

Q at equilibrium = 4

The producer surplus is

"4*(3+4^2)-\int(3+Q^2)dQ"

"19*4-(3Q+Q^3\/3)"

Since Q=4

"19*4-(3*4+4^3\/3)" =$42.60